1)American capitalism is not pure capitalism; it has become a "mixed" economy, largely because of defects or failures in the market system that Adam Smith did not foresee. What are these defects or failures that Smith didn't foresee? Discuss.
2) Income inequality in most European countries is smaller than it is in the US. Given all of the approaches and strategies available to the US (and to Europe) for improving income distribution, why do you think this is so?
1
Expert's answer
2018-03-05T08:34:07-0500
A) GDP = W + R + P + R = K60000 + K30000 + K10000 +K8000 = K108000
B) Level of saving in this economy = (K108000 – K101000) = K7000
C) Saving and investment are equal only in equilibrium; they are unequal under conditions of disequilibrium.
D) In a closed circular income stream, money flows continuously from firms to households. Households spend all of their money on goods and businesses spend all of their money on labor and expansion. The basic model of the circular flow of income ignores common consumer actions that take money out of the circular of income, or leakage. For instance, most households save money and some companies retain earnings.
E) The circular flow will adjust following new injections into it or new withdrawals from it. An injection of new spending will increase the flow. A net injection relates to the overall effect of injections in relation to withdrawals following a change in an economic variable.